# A Survey on Large Language Models for Code Generation

### v1

@@ 2.2. Code Generation
(13) Greedy $\textbf{Greedy Search}:\mathbf{y^{*}}=\mathop{\mathrm{argmax}}_{\mathbf{y}}P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k})),k=\{0,1,\dots,M\}$
(14) $\textbf{Sampling}:\mathbf{y}\sim P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y}_{i})\}_{i=1}^{k})),k=\{0,1,\dots,M\}$
@@ 3. Taxonomy
The recent surge in the development of Large Language Models (LLMs) has led to a significant number of these models being repurposed for code generation task through continued pre-training or fine-tuning. This trend is particularly observable in the realm of open-source models. For instance, Meta AI initially made the LLaMA (Touvron et al., 2023a) model publicly available, which was followed by the release of Code LLaMA (Roziere et al., 2023), designed specifically for code generation. Similarly, DeepSeek LLM (Bi et al., 2024a) developed and released by DeepSeeker has been extended to create DeepSeek Coder (Guo et al., 2024), a variant tailored for code generation. The Qwen team has developed and released Code Qwen (Team, 2024), building on their original Qwen (Bai et al., 2023) model. Microsoft, on the other hand, has unveiled WizardLM (Xu et al., 2023) and is exploring its coding-oriented counterpart, WizardCoder (Luo et al., 2023). Google has joined the fray by releasing Gemma (Team et al., 2024), subsequently followed by Code Gemma (CodeGemma Team et al., 2024). Beyond simply adapting general-purpose LLMs for code-related tasks, there has been a proliferation of models specifically engineered for code generation. Notable examples include StarCoder (Li et al., 2023a), OctoCoder (Muennighoff et al., 2023), and CodeGen (Nijkamp et al., 2022). These models underscore the trend of LLMs being developed with a focus on code generation.
Recognizing the importance of these developments, we propose a taxonomy that categorizes and evaluates the latest advances in LLMs for code generation. This taxonomy, depicted in Figure 3, serves as a comprehensive reference for researchers seeking to quickly familiarize themselves with the state-of-the-art in this dynamic field.
@@
Curation [Pre-training [CodeSearchNet(Husain et al., 2019), Google BigQuery(Hoffa, 2016), The Pile(Gao et al., 2020), CodeParrot(Tunstall et al., 2022), GitHub Code(Tunstall et al., 2022)
@@
Tuning [CommitPackFT (Muennighoff et al., 2023), Code Alpaca(Chaudhary, 2023), OA-Leet(Computations, 2023), OSS-Instruct(Wei et al., 2023), Evol-instruction(Roshdieh, 2023)
Self-OSS-Instruct-SC2-Exec-Filter(Yuxiang Wei, 2024) ,leaf,text width=24.5em] ] [Benchmarks [General [HumanEval(Chen et al., 2021), HumanEval+(Liu et al., 2024b), HumanEvalPack(Muennighoff et al., 2023), MBPP(Austin et al., 2021)
MBPP+(Liu et al., 2024b), CoNaLa(Yin et al., 2018), Spider(Yu et al., 2018), CONCODE(Iyer et al., 2018), ODEX(Wang et al., 2022d)
CoderEval(Yu et al., 2024), ReCode(Wang et al., 2022c), StudentEval(Babe et al., 2023) ,leaf,text width=19em] ] [Competitions [APPS(Hendrycks et al., 2021), CodeContests(Li et al., 2022a) ,leaf,text width=19em] ] [Data Science [DSP(Chandel et al., 2022), DS-1000(Lai et al., 2023), ExeDS(Huang et al., 2022) ,leaf,text width=19em] ] [Multilingual [MBXP(Athiwaratkun et al., 2022), Multilingual HumanEval(Athiwaratkun et al., 2022), HumanEval-X(Zheng et al., 2023c), MultiPL-E(Cassano et al., 2022)
xCodeEval(Khan et al., 2023) ,leaf,text width=19em] ] [Reasoning [MathQA-X(Athiwaratkun et al., 2022), MathQA-Python(Austin et al., 2021), GSM8K(Cobbe et al., 2021), GSM-HARD(Gao et al., 2023a) ,leaf,text width=19em] ] [Repository [RepoEval(Zhang et al., 2023c), Stack-Repo(Shrivastava et al., 2023a), Repobench(Liu et al., 2023b), EvoCodeBench(Li et al., 2024)
@@
Synthesis [Self-Instruct (Wang et al., 2023a), Evol-Instruct (Xu et al., 2023), Phi-1(Gunasekar et al., 2023), Code Alpaca(Chaudhary, 2023), WizardCoder(Luo et al., 2023)
Magicoder(Wei et al., 2023), StarCoder2-instruct (Yuxiang Wei, 2024) ,leaf,text width=24.5em] ] [Pre-training [Model
@@
OctoPack(Muennighoff et al., 2023), Code Llama(Roziere et al., 2023), MFTCoder(Liu et al., 2023a)
phi-1.5(Li et al., 2023b), CodeShell(Xie et al., 2024), Magicoder(Wei et al., 2023)
@@
StarCoder2-Instruct(Yuxiang Wei, 2024) ,leaf,text width=13em] ] ] [Pre-training
Tasks [CLM(Li et al., 2023a; Luo et al., 2023; Wei et al., 2023; Guo et al., 2024), DAE(Ahmad et al., 2021; Wang et al., 2021, 2023b), Auxiliary(Wang et al., 2021; Chai et al., 2022; Wang et al., 2023b) ,leaf,text width=16em] ] ] [Fine-tuning [Instruction
Tuning [Full Parameter
@@
CodeGeeX2(Zheng et al., 2023c), Magicoder(Wei et al., 2023), CodeGemma(CodeGemma Team et al., 2024)
@@
Fine-tuning [CodeUp(Jiang and Kim, 2023), ASTRAIOS(Zhuo et al., 2024) ,leaf,text width=8em] ] ] [Reinforcement
@@
with Feedback [CodeRL(Le et al., 2022), CompCoder(Wang et al., 2022a), PPOCoder(Shojaee et al., 2023), RLTF(Liu et al., 2023d)
@@
Engineering [Reflexion(Shinn et al., 2024), LATS(Zhou et al., 2023), Self-Debugging(Chen et al., 2023), SelfEvolve(Jiang et al., 2023)
Theo X. et al.(Olausson et al., 2023), CodeT(Chen et al., 2022b), LEVER(Ni et al., 2023a), AlphaCodium(Ridnik et al., 2024) ,leaf,text width=17em] ] [Repository
@@
Context [RepoCoder(Zhang et al., 2023c), CoCoMIC(Ding et al., 2022b), RepoHyper(Phan et al., 2024), RLPG(Shrivastava et al., 2023b)
Repoformer(Wu et al., 2024), RepoFusion(Shrivastava et al., 2023a), ToolGen(Wang et al., 2024c), CodePlan(Bairi et al., 2023)
@@
Augmented [HGNN(Liu et al., 2020), REDCODER(Parvez et al., 2021), ReACC(Lu et al., 2022), DocPrompting(Zhou et al., 2022a)
@@
Coding Agents [AgentCoder (Huang et al., 2023), MetaGPT(Hong et al., 2023), CodeAct (Wang et al., 2024a), AutoCodeRover (Zhang et al., 2024), Devin(Cognition, 2024)
OpenDevin(OpenDevin, 2024), SWE-agent(John Yang, 2024), L2MAC(Holt et al., 2023), OpenDevin CodeAct 1.0(Xingyao Wang and Neubig, 2024) ,leaf,text width=22em] ] ] [Evaluation [Metrics [Exact Match, BLEU(Papineni et al., 2002), ROUGE(Lin, 2004), METEOR(Banerjee and Lavie, 2005), CodeBLEU(Ren et al., 2020), pass@k(Chen et al., 2021)
@@
Evaluation [CodePlan(Bairi et al., 2023), RepoFusion(Shrivastava et al., 2023a), CodeBLEU(Ren et al., 2020) ,leaf,text width=13em] ] [LLM-as-a-Judge [AlpacaEval(Li et al., 2023c), MT-bench(Zheng et al., 2024a), ICE-Score(Zhuo, 2024) ,leaf,text width=13em] ] ] [Application [GitHub Copilot(Chen et al., 2021), CodeGeeX(Zheng et al., 2023c), CodeWhisperer(Amazon, 2022), Codeium(Codeium, 2023), CodeArts Snap(Shen et al., 2023), TabNine(TabNine, 2018)
Replit(Replit, 2016) ,leaf,text width=27.5em] ] ]
Figure 3. Taxonomy of large language models (LLMs) for code generation.

### v2

@@ 2.2. Code Generation
(13) Greedy $\textbf{Greedy Search}:\mathbf{y^{*}}=\mathop{\mathrm{argmax}}_{\mathbf{y}}P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k})),k\in\{0,1,\dots,M\}$
(14) $\textbf{Sampling}:\mathbf{y}\sim P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y}_{i})\}_{i=1}^{k})),k\in\{0,1,\dots,M\}$
To verify the functionality correctness of the generated code solution, $\mathbf{y}$ is subsequently executed via a compiler or interpreter, represented as $\mathbf{Exe}(\cdot)$, on a suit of unit tests $\mathbf{\mathcal{T}}$. The feedback from this execution can be denoted as $\mathbf{Feedback}(\mathbf{Exe}(\mathbf{y},\mathbf{\mathcal{T}}))$. If the generated code solution fails to pass all test cases, the error feedback can be iteratively utilized to refine the code by leveraging the previous attempt ($\mathbf{y}_{pre}$) and the associated feedback. Formally,
(15) $\mathbf{y}\sim P_{\theta}(\mathbf{y}\mid\operatorname{prompt}(\mathbf{x},\{(\mathbf{x_{i}},\mathbf{y_{i}})\}_{i=1}^{k},\mathbf{y}_{pre},\mathbf{Feedback}(\mathbf{Exe}(\mathbf{y},\mathbf{\mathcal{T}})))),k\in\{0,1,\dots,M\}$
Further details and relevant studies on using feedback to improve code generation are comprehensively discussed in Section 5.5 and 5.6.
3. Methodology
Figure 3. Overview of the paper search and collection process.
In this section, we detail the systematic methodologies employed in conducting literature reviews. We follow the systematic literature review methodology outlined by (Kitchenham et al., 2009), which has been widely adopted in numerous software engineering literature reviews (Hou et al., 2024; Li et al., 2017; Liu et al., 2022b; Ramirez et al., 2018; Wang et al., 2022a). The overall process is illustrated in Figure 3, and the detailed steps in our methodology are documented below.
3.1. Research Questions
To deliver a comprehensive and up-to-date literature review on the latest advancements in large language models (LLMs) for code generation, this systematic literature review addresses the following research questions (RQs):
RQ1: How can we categorize and evaluate the latest advances in LLMs for code generation? The recent proliferation of LLMs has resulted in many of these models being adapted for code generation task. While the adaptation of LLMs for code generation essentially follows the evolution of LLMs, this evolution encompasses a broad spectrum of research directions and advancements. For software engineering (SE) researchers, it can be challenging and time-consuming to fully grasp the comprehensive research landscape of LLMs and their adaptation to code generation. RQ1 aims to propose a taxonomy that serves as a comprehensive reference for researchers, enabling them to quickly familiarize themselves with the state-of-the-art in this dynamic field and identify specific research problems and directions of interest.
RQ2: What are the key insights into LLMs for code generation? RQ2 seeks to assist researchers in establishing a comprehensive, up-to-date, and advanced understanding of LLMs for code generation. This includes discussing various aspects of this rapidly evolving domain, such as data curation, latest advancements, performance evaluation, ethical and environmental implications, and real-world applications. A historical overview of the evolution of LLMs for code generation is provided, along with an empirical comparison using the widely recognized HumanEval and MBPP benchmarks, as well as the more practical and challenging BigCodeBench benchmark, to highlight the progressive enhancements in LLM capabilities for code generation. RQ2 offers an in-depth analysis of critical insights related to LLMs for code generation.
RQ3: What are the critical challenges and promising research opportunities in LLMs for code generation? Despite the revolutionary impact of LLMs on the paradigm of code generation and their remarkable performance, numerous challenges remain unaddressed. These challenges primarily stem from the gap between academic research and practical development. For instance, while the HumanEval benchmark is established as a de facto standard for evaluating the coding proficiency of LLMs in academia, it has been shown that this evaluation does not adequately reflect practical development scenarios (Jimenez et al., 2023; Du et al., 2024; Liu et al., 2024c; Ding et al., 2024). RQ3 aims to identify critical challenges and highlight promising opportunities to bridge the gap between research and practical application.
3.2. Search Process
Table 2. Publication venues for conference proceedings and journals articles for manual search.
Domain Venue Acronym
LLMs International Conference on Learning Representations ICLR
Conference on Neural Information Processing Systems NeurIPS
International Conference on Machine Learning ICML
Annual Meeting of the Association for Computational Linguistics ACL
Conference on Empirical Methods in Natural Language Processing EMNLP
International Joint Conference on Artificial Intelligence NAACL
AAAI Conference on Artificial Intelligence AAAI
SE International Conference on Software Engineering ICSE
Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering ESEC/FSE
International Conference on Automated Software Engineering ASE
Transactions on Software Engineering and Methodology TOSEM
Transactions on Software Engineering TSE
International Symposium on Software Testing and Analysis ISSTA
Table 3. Keywords related to LLMs and code generation task for automated search.
Keywords Related to LLMs Keywords Related to Code Generation Task
Code Large Language Model, Code LLMs, Code Language Model, Code LMs, Large Language Model, LLM, Language Model, LM, Pre-trained Language Model, PLM, Pre-trained model, Natural Language Processing, NLP, GPT-3, ChatGPT, GPT-4, LLaMA, CodeLlama, PaLM, CodeT5, Codex, CodeGen, InstructGPT Code Generation, Program Synthesis, Code Intelligence, Coder, natural-language-to-code, NL2Code, Programming
3.2.1. Search Strings
To address the aforementioned three research questions (RQs), we initiate a manual review of conference proceedings and journal articles from top-tier venues in the fields of LLMs and SE, as detailed in Table 2. This process allowed us to identify relevant studies and derive search strings, which are subsequently utilized for an automated search across various scientific databases. The complete set of search keywords is presented in Table 3.
3.2.2. Search Databases
Following the development of search strings, we executed an automated search using four popular scientific databases: the ACM Digital Library, IEEE Xplore Digital Library, arXiv, and DBLP. Our search focus on identifying papers whose titles contain keywords pertinent to LLMs and code generation. This approach enhances the likelihood of retrieving relevant papers since both sets of keywords must be present in the title. Although this title-based search strategy effectively retrieves a large volume of papers, it is important to note that in some instances (Shojaee et al., 2023), the scope of code generation can be broader, encompassing areas such as code completion, code translation, and program synthesis. As outlined in Section 1, this survey adopts a prevalent definition of code generation as the natural-language-to-code (NL2Code) task (Austin et al., 2021; Athiwaratkun et al., 2022; Zan et al., 2023).
Consequently, we conduct further automatic filtering based on the content of the papers. Papers focusing on “code completion” and “code translation” are excluded unless they pertain to the specific topics discussed in Section 5.7 and Section 5.8, where code completion is a primary focus. After completing the automated search, the results from each database are merged and deduplicated using scripts. This process yields 294 papers from arXiv, 73 papers from the ACM Digital Library, 36 papers from IEEE Xplore, and 261 papers from DBLP.
3.3. Inclusion and Exclusion Criteria
The search process conducted across various databases and venues is intentionally broad to gather a comprehensive pool of candidate papers. This approach maximizes the collection of potentially relevant studies. However, such inclusivity may lead to the inclusion of papers that do not align with the scope of this survey, as well as duplicate entries from multiple sources. To address this, we have established a clear set of inclusion and exclusion criteria, based on the guidelines from (Hou et al., 2024; Wang et al., 2024b). These criteria are applied to each paper to ensure alignment with our research scope and questions, and to eliminate irrelevant studies.
Inclusion Criteria. A paper will be included if it meets any of the following criteria:
  • It is available in full text.
  • It presents a dataset or benchmark specifically designed for code generation with LLMs.
  • It explores specific LLM techniques, such as pre-training or instruction tuning, for code generation.
  • It provides an empirical study or evaluation related to the use of LLMs for code generation.
  • It discusses the ethical considerations and environmental impact of deploying LLMs for code generation.
  • It proposes tools or applications powered by LLMs for code generation.
Exclusion Criteria. Conversely, papers will be excluded if they meet any of the following conditions:
  • They are not written in English.
  • They are found in books, theses, monographs, keynotes, panels, or venues (excluding arXiv) that do not undergo a full peer-review process.
  • They are duplicate papers or different versions of similar studies by the same authors.
  • They focus on text generation rather than source code generation, such as generating code comments, questions, test cases, or summarization.
  • They do not address the task of code generation, for instance, focusing on code translation instead.
  • They leverage software engineering methods to enhance code generation without emphasizing LLMs.
  • They do not utilize LLMs, opting for other models like Long Short-Term Memory (LSTM) networks.
  • They use encoder-only language models, such as BERT, which are not directly applicable to code generation task.
  • LLMs are mentioned only in future work or discussions without being central to the proposed approach.
Papers identified through both manual and automated searches undergo a detailed manual review to ensure they meet the inclusion criteria and do not fall under the exclusion criteria. Specifically, the first two authors independently review each paper to determine its eligibility. In cases of disagreement, the third author makes the final inclusion decision.
3.4. Quality Assessment
To ensure the inclusion of high-quality studies, we have developed a comprehensive set of ten Quality Assessment Criteria (QAC) following (Hou et al., 2024). These QAC are designed to evaluate the relevance, clarity, validity, and significance of the papers considered for our review.
In accordance with (Hou et al., 2024), the first three QAC assess the study’s alignment with our objectives. These criteria are rated as “irrelevant/unmet”, “partially relevant/met”, or “relevant/fully met”, corresponding to scores of -1, 0, and 1, respectively. If a study receive a score of -1 across these initial three criteria, it is deemed ineligible for further consideration and subsequently excluded from our review process.
The subsequent seven QAC focus on a more detailed content evaluation, employing a scoring range of -1 to 2, representing “poor”, “fair”, “good”, and “excellent”. We compute a cumulative score based on the responses to QAC4 through QAC10 for each paper. For published works, the maximum achievable score is 14 (2 points per question). We retain those with a score of 11.2 (80% of the total score) or higher. For unpublished papers available on arXiv, QAC4 defaults to a score of 0, making the maximum possible score for the remaining criteria 12. Accordingly, we retain papers scoring 9.6 (80% of the adjusted total score) or above .
  • QAC1: Is the research not classified as a secondary study, such as a systematic literature review or survey? (-1, 0, 1)
  • QAC2: Does the study incorporate the use of LLMs? (-1, 0, 1)
  • QAC3: Is the study relevant to the code generation task? (-1, 0, 1)
  • QAC4: Is the research published in a prestigious venue? (-1, 0, 1, 2)
  • QAC5: Does the study present a clear research motivation? (-1, 0, 1, 2)
  • QAC6: Are the key contributions and limitations of the study discussed? (-1, 0, 1, 2)
  • QAC7: Does the study contribute to the academic or industrial community? (-1, 0, 1, 2)
  • QAC8: Are the LLM techniques employed in the study clearly described? (-1, 0, 1, 2)
  • QAC9: Are the experimental setups, including experimental environments and dataset information, thoroughly detailed? (-1, 0, 1, 2)
  • QAC10: Does the study clearly confirm its experimental findings? (-1, 0, 1, 2)
3.5. Snowballing Search
Following the quality assessment, we establish an initial set of papers for our study. To minimize the risk of excluding pertinent literature, we implement a snowballing search strategy. Snowballing search involves utilizing a paper’s reference list or its citations to discover additional relevant studies, known as backward and forward snowballing, respectively. In this survey, we exclusively employed backward snowballing following (Wang et al., 2024b). Despite this effort, no additional studies are identified through this method. This could be attributed to the task-specific nature of the code generation (natural-language-to-code), where reference studies are typically published earlier. Consequently, our methodology, which encompassed an extensive manual and automated search, likely covered the relevant literature comprehensively, explaining the lack of additional studies through snowballing search.
Figure 4. Data qualitative analysis. Top: Annual distribution of selected papers across various publication venues. Bottom: Distribution analysis of research topics covered in the included papers.
3.6. Data Collection and Analysis
The data collection process for our study, illustrated in Figure 3, began with a manual search through conference proceedings and journal articles from leading venues in LLMs and SE. This initial step yielded 42 papers, from which we extracted relevant search strings. Following this, we performed an automated search across four academic databases using keyword-based queries, resulting in the retrieval of 664 papers. After performing automatic filtering (351 papers), applying inclusion and exclusion criteria (247 papers), conducting quality assessments (235 papers), and utilizing snowballing search (235 papers), we finalize a collection of 235 papers focusing on LLMs for code generation.
To provide insights from the selected papers, we begin by presenting an overview of their distribution across publication venues each year, as illustrated at the top of Figure 4. Our analysis indicates that 14% of the papers are published in LLM-specific venues and 7% in SE venues. Remarkably, 49% of the papers remain unpublished in peer-reviewed venues and are available on arXiv. This trend is understandable given the emerging nature of this field, with many works being recent and pending formal submission. Despite the absence of peer review on arXiv, our quality assessment process ensures that only high-quality papers are included, thereby maintaining the integrity of this survey. Furthermore, the annual trend in the number of collected papers indicates nearly exponential growth in the field. From a single paper in the period 2018 to 2020, the numbers increased to 6 in 2021, 11 in 2022, 75 in 2023, and 140 in 2024. This trend reflects growing interest and attention in this research area, with expectations for continued expansion in the future. Additionally, to capture the breadth of advancements in LLMs for code generation, we conducted a distribution analysis of the research topics covered in the included papers, as shown at the bottom of Figure 4. We observe that the development of LLMs for code generation closely aligns with broader trends in general-purpose LLM research. Notably, the most prevalent research topics are Pre-training and Foundation Models (21.5%), Prompting (11.8%), and Evaluation and Benchmarks (24.1%). These areas hold significant promise for enhancing, refining, and evaluating LLM-driven code generation.
@@ 4. Taxonomy
The recent surge in the development of LLMs has led to a significant number of these models being repurposed for code generation task through continual pre-training or fine-tuning. This trend is particularly observable in the realm of open-source models. For instance, Meta AI initially made the LLaMA (Touvron et al., 2023a) model publicly available, which was followed by the release of Code Llama (Roziere et al., 2023), designed specifically for code generation. Similarly, DeepSeek LLM (Bi et al., 2024a) developed and released by DeepSeek has been extended to create DeepSeek Coder (Guo et al., 2024), a variant tailored for code generation. The Qwen team has developed and released Code Qwen (Team, 2024), building on their original Qwen (Bai et al., 2023) model. Microsoft, on the other hand, has unveiled WizardLM (Xu et al., 2023) and is exploring its coding-oriented counterpart, WizardCoder (Luo et al., 2023). Google has joined the fray by releasing Gemma (Team et al., 2024), subsequently followed by Code Gemma (CodeGemma Team et al., 2024). Beyond simply adapting general-purpose LLMs for code-related tasks, there has been a proliferation of models specifically engineered for code generation. Notable examples include StarCoder (Li et al., 2023a), OctoCoder (Muennighoff et al., 2023), and CodeGen (Nijkamp et al., 2022). These models underscore the trend of LLMs being developed with a focus on code generation.
Recognizing the importance of these developments, we conduct a thorough analysis of selected papers on LLMs for code generation, sourced from widely used scientific databases as mentioned in Section 3. Based on this analysis, we propose a taxonomy that categorizes and evaluates the latest advancements in LLMs for code generation. This taxonomy, depicted in Figure 6, serves as a comprehensive reference for researchers seeking to quickly familiarize themselves with the state-of-the-art in this dynamic field. It is important to highlight that the category of recent advances emphasizes the core techniques used in the current state-of-the-art code LLMs.
@@
Figure 5. A diagram illustrating the general training, inference, and evaluation workflow for Code LLMs and their associated databases. The training workflow is mainly divided into four distinct stages: Stage ① and ② are the pre-training phase, whereas Stages ③ and ④ represent the post-training phases. It is important to note that Stage ② and ④ are optional. For instance, StarCoder (Li et al., 2023a) incorporates only Stage ①. WizardCoder (Luo et al., 2023), fine-tuned upon StarCoder, includes only Stage ③, while Code Llama (Roziere et al., 2023), continually pre-trained on Llama 2, encompasses Stages ② and ③. DeepSeek-Coder-V2 (Zhu et al., 2024), continually pre-trained on DeepSeek-V2, covers Stages ②, ③, and ④. Note that pre-trained model can be directly used for inference through prompt engineering.
@@
Curation
(Sec. 5.1) [Pre-training [CodeSearchNet(Husain et al., 2019), Google BigQuery(Hoffa, 2016), The Pile(Gao et al., 2020), CodeParrot(Tunstall et al., 2022), GitHub Code(Tunstall et al., 2022)
@@
Tuning [CommitPackFT (Muennighoff et al., 2023), Code Alpaca(Chaudhary, 2023), OA-Leet(Computations, 2023), OSS-Instruct(Wei et al., 2023b), Evol-instruction(Roshdieh, 2023)
Self-OSS-Instruct-SC2-Exec-Filter(Yuxiang Wei, 2024) ,leaf,text width=24.5em] ] [Benchmarks [General [HumanEval(Chen et al., 2021), HumanEval+(Liu et al., 2024c), HumanEvalPack(Muennighoff et al., 2023), MBPP(Austin et al., 2021)
MBPP+(Liu et al., 2024c), CoNaLa(Yin et al., 2018), Spider(Yu et al., 2018), CONCODE(Iyer et al., 2018), ODEX(Wang et al., 2022e)
CoderEval(Yu et al., 2024), ReCode(Wang et al., 2022d), StudentEval(Babe et al., 2023) ,leaf,text width=19em] ] [Competitions [APPS(Hendrycks et al., 2021), CodeContests(Li et al., 2022a) ,leaf,text width=19em] ] [Data Science [DSP(Chandel et al., 2022), DS-1000(Lai et al., 2023), ExeDS(Huang et al., 2022) ,leaf,text width=19em] ] [Multilingual [MBXP(Athiwaratkun et al., 2022), Multilingual HumanEval(Athiwaratkun et al., 2022), HumanEval-X(Zheng et al., 2023c), MultiPL-E(Cassano et al., 2022)
xCodeEval(Khan et al., 2023) ,leaf,text width=19em] ] [Reasoning [MathQA-X(Athiwaratkun et al., 2022), MathQA-Python(Austin et al., 2021), GSM8K(Cobbe et al., 2021), GSM-HARD(Gao et al., 2023a) ,leaf,text width=19em] ] [Repository [RepoEval(Zhang et al., 2023c), Stack-Repo(Shrivastava et al., 2023a), Repobench(Liu et al., 2023c), EvoCodeBench(Li et al., 2024)
@@
Synthesis
(Sec. 5.2) [Self-Instruct (Wang et al., 2023a), Evol-Instruct (Xu et al., 2023), Phi-1(Gunasekar et al., 2023), Code Alpaca(Chaudhary, 2023), WizardCoder(Luo et al., 2023)
Magicoder(Wei et al., 2023b), StarCoder2-instruct (Yuxiang Wei, 2024) ,leaf,text width=24.5em] ] [Pre-training
(Sec. 5.3) [Model
@@
OctoPack(Muennighoff et al., 2023), Code Llama(Roziere et al., 2023), MFTCoder(Liu et al., 2023b)
phi-1.5(Li et al., 2023b), CodeShell(Xie et al., 2024), Magicoder(Wei et al., 2023b)
@@
StarCoder2-Instruct(Yuxiang Wei, 2024), Codestral(MistralAI, 2024) ,leaf,text width=13em] ] ] [Pre-training
Tasks [CLM(Li et al., 2023a; Luo et al., 2023; Wei et al., 2023b; Guo et al., 2024), DAE(Ahmad et al., 2021; Wang et al., 2021, 2023b), Auxiliary(Wang et al., 2021; Chai et al., 2022; Wang et al., 2023b) ,leaf,text width=16.5em] ] ] [Fine-tuning [Instruction
Tuning
(Sec. 5.4) [Full Parameter
@@
CodeGeeX2(Zheng et al., 2023c), Magicoder(Wei et al., 2023b), CodeGemma(CodeGemma Team et al., 2024)
@@
Fine-tuning [CodeUp(Jiang and Kim, 2023), ASTRAIOS(Zhuo et al., 2024b) ,leaf,text width=8em] ] ] [Reinforcement
@@
with Feedback
(Sec. 5.5) [CodeRL(Le et al., 2022), CompCoder(Wang et al., 2022b), PPOCoder(Shojaee et al., 2023), RLTF(Liu et al., 2023e)
@@
Engineering
(Sec. 5.6) [Reflexion(Shinn et al., 2024), LATS(Zhou et al., 2023b), Self-Debugging(Chen et al., 2023), SelfEvolve(Jiang et al., 2023)
Theo X. et al.(Olausson et al., 2023), CodeT(Chen et al., 2022c), LEVER(Ni et al., 2023a), AlphaCodium(Ridnik et al., 2024) ,leaf,text width=17em] ] [Repository
@@
Context
(Sec. 5.7) [RepoCoder(Zhang et al., 2023c), CoCoMIC(Ding et al., 2022b), RepoHyper(Phan et al., 2024), RLPG(Shrivastava et al., 2023b)
Repoformer(Wu et al., 2024), RepoFusion(Shrivastava et al., 2023a), ToolGen(Wang et al., 2024d), CodePlan(Bairi et al., 2023)
@@
Augmented
(Sec. 5.8) [HGNN(Liu et al., 2020), REDCODER(Parvez et al., 2021), ReACC(Lu et al., 2022), DocPrompting(Zhou et al., 2022a)
@@
Coding Agents
(Sec. 5.9) [AgentCoder (Huang et al., 2023), MetaGPT(Hong et al., 2023), CodeAct (Wang et al., 2024a), AutoCodeRover (Zhang et al., 2024b), Devin(Cognition, 2024)
OpenDevin(OpenDevin, 2024), SWE-agent(John Yang, 2024), L2MAC(Holt et al., 2023), OpenDevin CodeAct 1.0(Xingyao Wang and Neubig, 2024) ,leaf,text width=22em] ] ] [Evaluation
(Sec. 5.10) [Metrics [Exact Match, BLEU(Papineni et al., 2002), ROUGE(Lin, 2004), METEOR(Banerjee and Lavie, 2005), CodeBLEU(Ren et al., 2020), pass@k(Chen et al., 2021)
@@
Evaluation [CodePlan(Bairi et al., 2023), RepoFusion(Shrivastava et al., 2023a), CodeBLEU(Ren et al., 2020) ,leaf,text width=13em] ] [LLM-as-a-Judge [AlpacaEval(Li et al., 2023c), MT-bench(Zheng et al., 2024a), ICE-Score(Zhuo, 2024) ,leaf,text width=13em] ] ] [Code LLMs
Alignment
(Sec. 5.11) [Green(Shi et al., 2024; Wei et al., 2023a), Responsibility(Liu et al., 2023a; Xu et al., 2024b), Efficiency(Yang et al., 2024a), Safety(Schuster et al., 2021; Hajipour et al., 2024; Yang et al., 2024b; Al-Kaswan et al., 2024; Yuan et al., 2023a; Fried et al., 2022; Allal et al., 2023), Trustworthiness(Ji et al., 2023a; Palacio et al., 2023) ,leaf,text width=29.8em] ] [Application
(Sec. 5.12) [GitHub Copilot(Chen et al., 2021), CodeGeeX(Zheng et al., 2023c), CodeWhisperer(Amazon, 2022), Codeium(Codeium, 2023), CodeArts Snap(Shen et al., 2023), TabNine(TabNine, 2018), Replit(Replit, 2016) ,leaf,text width=29.8em] ] ]
Figure 6. Taxonomy of LLMs for code generation.